# Enhancing Quantum Memory Lifetime with Measurement-Free Local Error Correction and Reinforcement Learning

### v1

@@ III.2 Simplest LEC operation and its limitation
In contrast, as shown in the left of Fig. 4(b) and (c), the 2D Ising model and 4D toric code have 2D membranes of data qubit errors with 1D boundaries of syndromes. For these codes, repeating Toom’s rule operation is an established cellular automata decoding protocol [74, 75]. This operation can be implemented with CCX or CCZ gate as depicted in Fig. 4(b) and effectively “shrinks” the size of a 2D error membrane. In the limit where a system size $L\to\infty$, the repetition of this local operation followed by syndrome extraction operation can remove all errors in the system if the error density of data qubits is below some threshold [78, 80, 72]. Due to this property, the 2D Ising model and 4D toric code are called self-correcting memories.
However, for finite-sized systems, we observe error patterns that are uncorrectable by the simplest error removal operation as shown in Fig. 4. These error patterns limit the performance of the local decoding protocol for each code. In 2D toric code, the simplest error removal operation cannot remove nor reduce the error chains with length $\geq 2$. Also, for the finite-sized 2D Ising model and 4D toric code, some error configurations are uncorrectable by repeated Toom’s rule operations — even with perfect gates in the system below the threshold error density of data qubits. We illustrate the examples of such error patterns uncorrectable by repeated Toom’s rule operation as in Fig. 4(b) and (c) for the system size $L=4$. These configurations form periodic 2D sheets of data qubit errors surrounded by parallel 1D loops of syndromes. We call these error patterns error sheets with width defined as the distance between the boundary 1D syndrome loops. These uncorrectable errors were reported as energy barriers in previous literature [81, 44, 82]. Therefore, each error-correcting code has an issue of error patterns uncorrectable by the simplest error removal operation arising from different reasons: the operation’s limitation for 2D toric code and a system’s finiteness for 2D Ising model and 4D toric code.
III.3 Extended LEC operations
Now, we introduce how to resolve this issue of uncorrectable error patterns by adding error removal operations. We design a set of operations for each code such that any single low-weight error pattern can, in principle, be corrected.
@@ III.3 Extended LEC operations
For 2D toric code, we categorize the simplest removal operations as $d=1$ operations, because they can reduce length-1 error chains into length-0 error chains, i.e. no error. As shown in Fig. 5(a), we generalize this to $d=N$ operations that can reduce length-$N$ error chains into length-$(N-1)$ error chains. We include up to $d\leq 3$ operations in our optimization, where details are discussed in Appendix A.
In the 2D Ising model and 4D toric code, we extend the simplest operations in two ways as shown in Fig. 5(b). First, we include the Toom Sweep operations generated by 90 rotations of the Toom operations, which were shown to reduce the occurrence of uncorrectable membrane-like error patterns [40, 41]. Moreover, to enable the removal of such error patterns, we additionally introduce operations designed to remove errors between the two parallel 1D boundaries of syndromes. These operations take the same form as the $d=N$ operations for 2D toric code and can reduce width-$N$ error sheets to have narrower widths. We include two lowest-weight operations for both codes —- $d=1$ and $d=2$ operations for the 2D Ising model and $d=1$ and $d=(1,1)$ operations for the 4D toric code (see Appendix A for details).
@@ IV RL Optimization Framework
IV.1 RL Setting: Agent and Environment
RL is a machine learning paradigm where an agent learns which actions to play to maximize the reward through its interplay with an environment [52]. We consider the RL training scheme where the LEC circuit builder (or agent) learns which sequence of LEC gate operations (or actions) maximizes the decoding success rate (or reward) through its interplay with the circuit-level simulator (or environment) until the convergence of the agent-built LEC circuit. Due to the absence of mid-circuit readout in our scheme, our environment does not provide the observation to the agent, unlike traditional RL setting [52, 83]. This scenario is often called an open-loop control problem in the RL literature [84, 85, 86].
Figure 6: Overview of our reinforcement learning (RL) framework that consists of an agent, environment, and their interactions to optimize the LEC circuit for a given code and error model. By choosing the $(h+1)$-th action based on all previous actions, the agent’s policy network builds an LEC circuit, which is a sequence of fixed circuit depth $H$ number of LEC actions, as the environment’s input. Then, the environment benchmarks a decoding success rate given this LEC circuit with a fixed error model by simultaneously evaluating the success or failure of decoding for $N$ copies of codes after multiple LEC cycles, which are applications of LEC circuits after the introduction of new errors. This rate is then used as a reward to update the agent’s neural network parameters. These steps are repeated until the perceived reward saturates, i.e. until the agent receives the same reward signal and does not design a new LEC circuit for multiple network updates (see Appendix B.1 for details).
The specific interplay between the RL agent and environment is as summarized in Fig. 6 and Algorithm 1. First, the RL agent produces an LEC circuit based on its internal parameters. Then, the RL environment takes this LEC circuit as an input to compute its decoding success rate with fixed code and error model parameters. Such rate is used as a reward to update the internal parameters of the RL agent. This training process is repeated to maximize this reward until the resulting LEC circuit converges.
@@ IV.1 RL Setting: Agent and Environment
Let’s further discuss each component of our RL framework: an agent, environment, and algorithm updating the agent’s neural network based on reward from the environment.
First, the agent has a neural network $f^{\vec{\theta}}$ with vectorized internal parameters $\vec{\theta}$ that takes a sequence of $h$ (for $h<H$) actions as input and returns another action as output, i.e.
@@
where $\mathcal{A}$ denotes a set of all action candidates that depend on a choice of the error-correcting code. Repeating such next-action decision-making for $H$ times results in building the LEC circuit with the circuit depth $H$. Note that the size of the action space $|\mathcal{A}|$ depends on the code: $|\mathcal{A}|=24$ for 2D toric code, $|\mathcal{A}|=12$ for 2D Ising model, and $|\mathcal{A}|=26$ for 4D toric code (see Appendix A for details). Also, in practice, we use two fully connected hidden layers with 128 nodes each for both policy and value networks. We further discuss the details for the RL agent in Appendix B.2 and the effect of neural network size on RL training results in Appendix B.3.
Furthermore, the RL environment computes the decoding success rate by a circuit-level Monte Carlo simulation as illustrated in Fig. 6(b). This simulation takes a code type, error model parameters, and LEC circuit as inputs. Then, its reward computation is done by evaluating the average performance of an LEC circuit produced by the RL agent, because the decoder cannot access syndrome information in our measurement-free scheme.
We first run a simulation that applies multiple rounds of LEC. In each round, errors are introduced to data qubits, and then the input LEC circuit (fixed across all rounds) is applied to data qubits and newly initialized ancillas. Note that we are interested in extending the lifetime of memory encoded in the error-correcting code, and local decoding is performed multiple times through the lifetime of memory. Thus, we aim to reduce the accumulated error density over multiple rounds. We consider two different error sources:
  • Ambient errors introduced to all data qubits every LEC round before applying LEC circuit with error rate $p_{err}=p_{amb}$,
  • Gate errors introduced to all data qubits and ancillas after each parallelized multi-qubit gate operation with error rate $p_{err}=p_{gate}$.
We adopt an unbiased error model with bit-flip and dephasing errors. For 2D and 4D toric code, we apply $\hat{X}$ and $\hat{Z}$ error independently to each qubit with the same probability of $p_{err}$ ($\hat{Y}\sim\hat{X}\hat{Z}$ error applied with a probability of $p_{err}^{2}$). For the 2D Ising model, we apply $\hat{X}$ error independently to each qubit with the probability of $p_{err}$.
After multiple LEC rounds, we evaluate the decoding success. This final recovery step is to determine whether the final state of data qubits after multiple LEC rounds is “close enough” to the trivial ground state of the error-correcting code. This method varies among the choice of the code and is elaborated in Appendix B.4. We repeat this simulation for $N$ times with a fixed LEC circuit. For each simulation, a stochastic error model results in different configurations of errors applied to qubits. Thus, we can obtain the decoding success rate through this benchmark simulation.
Lastly, for updating the agent’s network parameters based on the reward from the environment, we adopt a proximal policy optimization (PPO) algorithm [83]. This algorithm has been widely used for optimizing a sequence of actions to perform a difficult task [87, 88]. We utilize the Python-based package Stable-Baselines3 to perform this training algorithm to implement the RL framework [89]. Note that the PPO algorithm has several hyperparameters for its training. We discuss the trials and observations on tuning the hyperparameter set to optimize the RL training result in Appendix B.3. To obtain the best performing RL-optimized LEC circuit, we run the RL training four times independently until the training reaches its termination condition (see Appendix B.1). Then, among the final LEC circuits from four trained models, we choose the one that maximizes the reward.
@@ IV.2 Optimizing depth and layout of LEC circuit
Figure 7: Optimization of depth and layout of the LEC circuit by the RL framework in each error-correcting code. We use code and error model parameters as in Table 1. Each final reward is computed with 10,000 independent samples. Our computation for the error bar of the final reward is discussed in Appendix C. (a) For 2D toric code, we show the RL training with variable depth with training $p_{gate}=2\times 10^{-4}$: as number of neural network (NN) policy updates increases, the LEC circuit’s variable depth converges. By comparing the final reward of the resulting LEC circuit trained by the RL framework with fixed vs. variable depth setting, we confirm that the obtained circuit depth is effectively optimized. (b) For the 2D Ising model and 4D toric code, we show the final reward after RL training vs. fixed circuit depth with training $p_{gate}=3\times 10^{-3}$ for 2D Ising model and $2\times 10^{-5}$ for 4D toric code. (c) For each training $p_{gate}$, we classify the components of RL-optimized LEC circuits as categorized in Fig. 5.
$L$ $p_{amb}$ $N$ $R$
2D toric code 8 0.02 100 5
2D Ising model 8 0.40 100 1
4D toric code 4 0.03 50 2
Table 1: Parameters for error-correcting code and for error model used for optimizing the LEC circuit with our RL framework. Note that $L$ denotes the system size, $p_{amb}$ denotes the ambient error rate in each LEC cycle before applying the LEC circuit, $N$ denotes the number of copies of code, and $R$ denotes the number of repeated LEC cycles used to compute decoding success rate as illustrated in Fig. 6. We discuss the reasoning why we choose these parameters in Appendix B.4.
@@
In each error-correcting code, we provide varying gate error rates $p_{gate}$ as an input to our RL framework to optimize the length and layout of LEC circuits for each input $p_{gate}$. We choose the scale of $p_{gate}$ to be $\sim 10^{-4}$ for 2D toric code, $\sim 10^{-3}$ for 2D Ising model, and $\sim 10^{-5}$ for 4D toric code. Then, we fix the other parameters for the code and the error model as shown in Table 1 to perform RL training.
@@
From RL training with varying fixed circuit depth $H$, we observe that there exists an optimal circuit depth for 2D toric code as shown in Fig. 7(a). Thus, we construct the RL framework that optimizes both circuit depth and layout (see Appendix B.2 for details). As shown in Fig. 7(c), we observe that the optimized circuit depth $H_{opt}$ for 2D toric code depends on training $p_{gate}$: $H_{opt}$ decreases as training $p_{gate}$ increases. This trend is because faulty LEC action causes a trade-off between error removal by parallel CCX or CCZ gate operation and error introduced by a gate error model. Even in $p_{gate}=0$ case, error removal by $h$-th LEC action $A_{h}$ gets marginalized as $h$ increases. When $p_{gate}$ increases, $A_{h}$ for large $h$ introduces more errors than it removes. We further observe that higher-weight LEC actions ($d>1$) are more frequently used for smaller training $p_{gate}$. This trend is because higher-weight error chains are less common than lower-weight error chains in the lattice and thus lower-weight LEC actions benefit more from the tradeoff of faulty LEC actions for smaller training $p_{gate}$.
@@
Unlike in 2D toric code, $H>H_{c}\approx 45$ shows a saturation of final reward in the self-correcting memories as shown in Fig. 7(b). This is consistent with the system entering a steady state, where an action coming after some $H_{c}$ achieves a balance between errors that it newly introduces by finite gate error or removes. Note that as $H\to\infty$, we expect the reward to slowly go to zero due to the rare possibility of high-weight uncorrectable errors. Thus, for the 2D Ising model and 4D toric code, we use the RL framework that optimizes circuit layout with fixed $H=60$ (see Appendix B.2 for details). However, just as in 2D toric code, Fig. 7(c) also shows that new LEC actions ($d=N$) are more frequently used for smaller training $p_{gate}$ with the same reasoning. Moreover, in both the 2D Ising model and 4D toric code, these new LEC actions appear only after a repetition of multiple Toom Sweep actions. For all training $p_{gate}$, RL-optimized LEC circuits start with $\geq 10$ Toom Sweep actions for 2D Ising model and $\geq 19$ Toom Sweep actions for 4D toric code. This is because the membrane-like uncorrectable error configurations illustrated in Fig. 4 appear only after applying an iteration of Toom’s rule decoding.
@@ V Memory Lifetime Extension by LEC
In this section, we provide a thorough analysis of the memory lifetime extension by the LEC circuits and the comparison between RL-optimized LEC circuits (or RL LEC) vs. conventional LEC circuits (or conventional LEC) in each code. Section V.1 analyzes the suppression of uncorrectable errors by RL LEC in comparison with conventional LEC. Section V.2 explains the idea of effective code distance that characterizes the lifetime of a memory encoded in the error-correcting code. Then, Section V.3 provides a comparison of this parameter for RL vs. conventional LEC based on the benchmark simulation results.
V.1 Effect of LEC circuits on error statistics
Figure 8: Classification of remaining errors after LEC cycle(s) in each code. Let a single LEC action be any LEC action among extended error removal operations as in Fig. 5. Then, we perform a benchmark simulation for applying LEC cycle(s) using code and error model parameters as in Table 1 with 10,000 independent samples. Note that for each LEC circuit, $p_{gate}$ is $1\times 10^{-4}$ for 2D toric code, $1\times 10^{-3}$ for 2D Ising model, and $1\times 10^{-5}$ for 4D toric code. Each error pattern after LEC cycle(s) is either (1) a non-logical low-weight error pattern correctable by one of single LEC actions, (2) a non-logical high-weight error pattern uncorrectable by any single LEC action, or (3) a logical error undetectable by extracting syndromes. We compare the occurrence of each category between RL-optimized LEC circuits (blue) vs. conventional LEC circuits in each code (grey). More detailed classification is presented in Appendix D.
In Section IV.2, we confirm that the RL LEC performs better than the conventional LEC in each error-correcting code. Recall that the conventional LEC for 2D toric code is called a nearest-neighbor LEC circuit (or NN LEC[46] and that for 2D Ising model and 4D toric code is called a Toom’s rule-based LEC circuit (or Toom LEC[38]. More specifically, RL LEC shows a lower decoding failure rate compared to conventional LEC after a certain number of repeated LEC cycles. To analyze the reason for this behavior, we classify remaining error patterns after repeated LEC cycle(s) in each code as shown in Fig. 8. Although both RL and Toom LEC show similar occurrences of logical errors, we observe that RL LEC is effective in suppressing non-logical errors that are either correctable or uncorrectable by a single LEC action that consists of the circuits. We discuss more details about this classification of remaining errors after repeated LEC cycles in Appendix D.

### v2

@@ III.2 Extended LEC operations
In contrast, as shown in the left of Fig. 4(b) and (c), the 2D Ising model and 4D toric code have 2D membranes of data qubit errors with 1D boundaries of syndromes. For these codes, repeating Toom’s rule operation is an established cellular automata decoding protocol [74, 75]. This operation can be implemented with CCX or CCZ gate as depicted in Fig. 4(b) followed by syndrome extraction. As this operation effectively “shrinks” the size of a 2D error membrane, its repetition can remove all errors in an infinite-sized system if the error density of data qubits is below some threshold [78, 80, 72]. Due to this property, the 2D Ising model and 4D toric code are called self-correcting memories.
However, we observe error patterns that are uncorrectable by the simplest error removal operation as shown in Fig. 4. In 2D toric code, the simplest error removal operation cannot remove nor reduce the error chains with length $\geq 2$. Also, for the finite-sized 2D Ising model and 4D toric code, some error configurations are uncorrectable by repeated Toom’s rule operations — even if the error density of data qubits is below the threshold. These configurations, also called as energy barriers [81, 44, 82], form periodic 2D sheets of data qubit errors surrounded by parallel 1D loops of syndromes. We call these error patterns error sheets with width defined as the distance between the boundary 1D syndrome loops. Therefore, each error-correcting code has an issue of error patterns uncorrectable by the simplest error removal operation arising from different reasons: the operation’s limitation for 2D toric code and a system’s finiteness for 2D Ising model and 4D toric code.
Now, we introduce how to resolve this issue of uncorrectable error patterns that limit the performance of the local decoding protocol. We design a set of operations for each code such that any single low-weight error pattern can, in principle, be corrected.
@@
For 2D toric code, we categorize the simplest removal operations as $d=1$ operations, because they can reduce length-1 error chains into length-0 error chains, i.e. no error. As shown in Fig. 5(a), we generalize this to $d=N$ operations that can reduce length-$N$ error chains into length-$(N-1)$ error chains. We include up to $d\leq 3$ operations in our optimization.
In the 2D Ising model and 4D toric code, we extend the simplest operations in two ways as shown in Fig. 5(b). First, we include the Toom Sweep operations generated by 90 rotations of the Toom operations, which were shown to reduce the occurrence of uncorrectable membrane-like error patterns [40, 41]. Moreover, to enable the removal of such error patterns, we additionally introduce operations designed to remove errors between the two parallel 1D boundaries of syndromes. These operations take the same form as the $d=N$ operations for 2D toric code and can reduce width-$N$ error sheets to have narrower widths. We include two lowest-weight operations for both codes — $d=1$ and $d=2$ operations for the 2D Ising model and $d=1$ and $d=(1,1)$ operations for the 4D toric code.
@@ IV RL Optimization Framework
RL is a machine learning paradigm where an agent learns which actions to play to maximize the reward through its interplay with an environment [52, 83]. In our setting, the LEC circuit builder (or agent) learns which sequence of LEC gate operations (or actions) maximizes the decoding success rate (or reward) through its interplay with the circuit-level simulator (or environment) until the convergence of the agent-built LEC circuit. Due to the absence of mid-circuit readout in our scheme, our environment does not provide the observation to the agent, unlike traditional RL setting [52, 83].
We aim to optimize an ordered sequence of discrete actions where an optimizer only observes the reward of the full sequence. Since the sequence is not adjusted during the execution and the reward structure is “opaque” to the optimizer, our setting can be classified as an open-loop control or black-box sequential optimization. Our problem features (1) a large search space of size between $10^{55}$ and $10^{85}$ possible ordered sequences, (2) the importance of discrete actions’ order in the sequence, and (3) costly reward computation — especially for 4D toric code.
Due to these features, we choose a proximal policy optimization (PPO) among RL techniques as our optimizer [84]. The RL algorithms have been used to robustly and sample-efficiently optimize the sequence of actions in large discrete action spaces [85, 86, 87, 88]. In particular, such algorithms have been successfully applied to the open-loop control problems [89, 90, 91]. More recently, the PPO algorithm has been used to optimize the quantum gate sequence, such as preparing multi-qubit GHZ states [92] or fault-tolerant logical states in small quantum error-correcting codes [70]. Although there are alternative optimizing strategies including variations of genetic algorithms [93] or simulated annealing [94], we find that the RL method is particularly suitable considering the purpose and features of our problem.
Figure 6: Overview of our reinforcement learning (RL) framework that consists of an agent, environment, and their interactions to optimize the LEC circuit for a given code and error model. By choosing the $(h+1)$-th action based on all previous actions, the agent’s policy network builds an LEC circuit, which is a sequence of fixed circuit depth $H$ number of LEC actions, as the environment’s input. Then, the environment benchmarks a decoding success rate given this LEC circuit with a fixed error model by simultaneously evaluating the success or failure of decoding for $N$ copies of codes after multiple LEC cycles, which are applications of LEC circuits after the introduction of new errors. This rate is then used as a reward to update the agent’s neural network parameters. These steps are repeated until the perceived reward saturates, i.e. until the agent receives the same reward signal and does not design a new LEC circuit for multiple network updates.
IV.1 Setting of RL framework
We summarize the specific interplay between the RL agent and environment as in Fig. 6 and Algorithm 1.
@@ IV.1 Setting of RL framework
First, the RL agent produces an LEC circuit based on its internal parameters. Then, the RL environment takes this LEC circuit as an input to compute its decoding success rate with fixed code and error model parameters. Such a rate is used as a reward to update the internal parameters of the RL agent. This training process is repeated to maximize this reward until the resulting LEC circuit converges.
Let’s further discuss each component of our RL framework: an agent, environment, and algorithm updating the agent’s neural network based on reward from the environment:
  • The agent has a neural network $f^{\vec{\theta}}$ with vectorized internal parameters $\vec{\theta}$ that takes a sequence of $h$ (for $h<H$) actions as input and returns the next action as output, i.e.
@@
    where $\mathcal{A}$ denotes a set of all action candidates that depend on a choice of the error-correcting code. Repeating such next-action decision-making for $H$ times results in building the LEC circuit with the circuit depth $H$. See Appendix C.1 for more details.
  • The environment computes the decoding success rate by a circuit-level Monte Carlo simulation of performing multiple rounds of applying an LEC circuit produced by the RL agent. This simulation takes a code type, error model parameters, and LEC circuit as inputs. Here, we consider two different error sources: (1) ambient errors introduced every LEC round before applying the LEC circuit with the rate $p_{\rm{amb}}$ and (2) gate errors introduced after applying each parallelized multi-qubit gate with rate $p_{\rm{gate}}$. Note that the reward computation is based on the average performance of an LEC circuit produced by the RL agent, because the decoder cannot access syndrome information in our measurement-free scheme. See Appendix C.2 for more details.
  • We utilize the Python-based package Stable-Baselines3 to perform the training to implement the RL framework [95]. Recall that we update the agent’s network parameters from the reward with the PPO algorithm. To obtain the best-performing LEC circuit, we run the RL training four times independently until the training reaches its termination condition. Among the final LEC circuits from four trained models, we choose the one that maximizes the reward.
Appendix C includes more details on our RL framework: a training termination condition (see Appendix C.3), hyperparameters (see Appendix C.4), and stability of training (see Appendix C.5).
@@ IV.2 Optimizing depth and layout of LEC circuit
In each error-correcting code, we provide varying gate error rates $p_{\rm{gate}}$ as an input to our RL framework to optimize the length and layout of LEC circuits for each input $p_{\rm{gate}}$. We choose the scale of $p_{\rm{gate}}$ to be $\sim 10^{-4}$ for 2D toric code, $\sim 10^{-3}$ for 2D Ising model, and $\sim 10^{-5}$ for 4D toric code (see Table 4 in Appendix C.2 for other code and error model parameters).
Figure 7: Optimization of depth and layout of the LEC circuit by the RL framework in each error-correcting code. Each final reward is computed with 10000 independent samples (see Appendix D for the error bar). (a) For 2D toric code, we show the RL training with variable depth with training $p_{\rm{gate}}=2\times 10^{-4}$: as number of neural network (NN) policy updates increases, the LEC circuit’s variable depth converges. By comparing the final reward of the resulting LEC circuit trained by the RL framework with fixed vs. variable depth setting, we confirm that the obtained circuit depth is effectively optimized. (b) For the 2D Ising model and 4D toric code, we show the final reward after RL training vs. fixed circuit depth with training $p_{\rm{gate}}=3\times 10^{-3}$ for 2D Ising model and $2\times 10^{-5}$ for 4D toric code. (c) For each training $p_{\rm{gate}}$, we classify the components of RL-optimized LEC circuits as categorized in Fig. 5.
@@
From RL training with varying circuit depth $H$, we observe that there exists an optimal circuit depth for 2D toric code as shown in Fig. 7(a). As shown in Fig. 7(c), we observe that the optimized circuit depth $H_{\rm{opt}}$ for 2D toric code depends on training $p_{\rm{gate}}$: $H_{\rm{opt}}$ decreases as training $p_{\rm{gate}}$ increases. This trend is because faulty LEC action causes a trade-off between error removal by parallel CCX or CCZ gate operation and error introduced by a gate error model. Even in $p_{\rm{gate}}=0$ case, error removal by $h$-th LEC action $A_{h}$ gets marginalized as $h$ increases. When $p_{\rm{gate}}$ increases, $A_{h}$ for large $h$ introduces more errors than it removes. We further observe that higher-weight LEC actions ($d>1$) are more frequently used for smaller training $p_{\rm{gate}}$. This trend is because higher-weight error chains are less common than lower-weight error chains in the lattice and thus lower-weight LEC actions benefit more from the tradeoff of faulty LEC actions for smaller training $p_{\rm{gate}}$.
@@
Unlike in 2D toric code, $H>H_{\rm{c}}\approx 45$ shows a saturation of final reward in the self-correcting memories as shown in Fig. 7(b). This is consistent with the system entering a steady state, where an action coming after some $H_{\rm{c}}$ achieves a balance between errors that it newly introduces by finite gate error or removes. Note that as $H\to\infty$, we expect the reward to slowly go to zero due to the rare possibility of high-weight uncorrectable errors. Thus, for the 2D Ising model and 4D toric code, we use the RL framework that optimizes circuit layout with fixed $H=60$ (see Appendix C.1 for details). However, just as in 2D toric code, Fig. 7(c) also shows that new LEC actions ($d=N$) are more frequently used for smaller training $p_{\rm{gate}}$ with the same reasoning. Moreover, in both the 2D Ising model and 4D toric code, these new LEC actions appear only after a repetition of multiple Toom Sweep actions. For all training $p_{\rm{gate}}$, RL-optimized LEC circuits start with $\geq 10$ Toom Sweep actions for 2D Ising model and $\geq 19$ Toom Sweep actions for 4D toric code. This is because the membrane-like uncorrectable error configurations illustrated in Fig. 4 appear only after applying an iteration of Toom’s rule decoding.
@@ V Memory Lifetime Extension by LEC
In this section, we provide a thorough analysis of the memory lifetime extension by the LEC circuits and the comparison between RL-optimized LEC circuits (or RL LEC) vs. conventional LEC circuits (or conventional LEC) in each code. Recall that the conventional LEC for 2D toric code is called a nearest-neighbor LEC circuit (or NN LEC[46] and that for 2D Ising model and 4D toric code is called a Toom’s rule-based LEC circuit (or Toom LEC[38]. Section V.1 explains the idea of effective code distance that characterizes the lifetime of a memory encoded in the error-correcting code. Then, Section V.2 provides a comparison of this parameter for RL vs. conventional LEC based on the benchmark simulation results. To explain why RL LEC performs better than conventional LEC, we provide the error statistics analysis in Appendix E.1.
V.1 Idea of effective code distance